How do you solve this equation OF Linear equations by substitution :
y=3-2x y=2-3x

Answer by rapaljer(4671)

Since y=3-2x, and y=2-3x, you can substitute the y = of the first equation into the y = of the second equation, and you have:
3-2x= 2-3x

Add +3x to each side:
3-2x + 3x = 2 -3x + 3x
3+x = 2

Subtract 3 from each side:
x=-1

Substitute back into the first equation:
y = 3-2x
y = 3-2(-1)
y = 5

Check in the second equation:
y = 2-3x
5 = 2-3(-1)
5=5
It checks!

The solution is the point (-1,5).

R^2 at SCC